# FunAudio-ASR Technical Report

### v1

@@ 1 Introduction
  • Optimization for practical production usage. Beyond achieving state-of-the-art performance on standardized benchmarks, FunAudio-ASR is meticulously engineered to meet the complex demands of real-world deployment scenarios, with a particular focus on practical usability, reliability, and user experience. We implement a comprehensive suite of optimizations in multiple dimensions, each addressing specific challenges encountered in commercial applications. (1) First, we implement a highly efficient streaming ASR architecture for FunAudio-ASR that supports real-time processing with minimal latency, enabling seamless integration into live applications such as video conferencing, live captioning, and voice-controlled devices. (2) Second, we enhance noise robustness substantially through a multi-stage approach. (3) Third, we implement advanced code-switching capabilities that seamlessly handle transitions between Chinese and English within the same utterance, which is critical for multilingual users in global business environments. (4) Fourth, we integrate customizable hotword recognition that allows users to define domain-specific terms or phrases for enhanced recognition accuracy. This feature is particularly valuable in specialized domains such as healthcare, enterprise, and automotive technology. Hotword recognition achieves
    To thoroughly evaluate these production-oriented optimizations, we develop a comprehensive evaluation protocol that includes both standardized benchmarks and real-world usage scenarios. This protocol encompasses distinct test sets, each simulating different application contexts. Our evaluation results reveal that FunAudio-ASR not only excels in recognition accuracy but also provides superior practical performance. These results demonstrate that FunAudio-ASR successfully bridges the gap between academic research and commercial production readiness, offering a comprehensive solution for addressing real-world speech recognition challenges.
@@ 2 Model Architecture
FunAudio-ASR comprises four key components: (1) an audio encoder that extracts representations from the input speech, implemented as multiple layers of transformer encoder, (2) an audio adaptor that connects the audio encoder output with the LLM, implemented as two layers of transformer encoder, (3) a CTC decoder that is built upon the audio encoder to obtain the initial recognition hypothesis, which will be used for hotword customization, as described in Section 5.5, and (4) an LLM-based decoder that produces output based on the audio condition and CTC prediction.
To address varying computational resource constraints and inference efficiency requirements, we propose two models with different sizes: FunAudio-ASR and FunAudio-ASR-nano. FunAudio-ASR comprises a 0.7B encoder and a 7B LLM-based decoder, aiming for the highest recognition accuracy, and FunAudio-ASR-nano comprises a 0.2B encoder and a 0.6B LLM-based decoder, seeking to strike a balance between accuracy and efficiency to meet the demands of low-resource scenarios.
@@ 3.2 Supervised Fine-tuning Data
The supervised fine-tuning (SFT) data consist of approximately millions of hours of data, including: human-transcribed data, pseudo-labeled data, environmental noise data, CosyVoice3 (Du et al., 2025) TTS generated data, simulated streaming data, noise augmented data and hotword customized data.
@@ 4.1 Pre-training of Audio Encoder
To develop a robust and effective audio encoder for integration into an LLM-based ASR (LLM-ASR) system, we adopt two complementary approaches, as shown in Figure 3. These strategies aim to leverage both self-supervised and supervised learning paradigms to produce high-quality speech representations that can be effectively aligned with linguistic knowledge in the LLM.
Stage 1: Self-supervised pre-training via Best-RQ framework with pre-trained text LLM initialization. The first stage leverages self-supervised pretraining through the Best-RQ (BERT-based Speech pre-Training with Random-projection Quantizer) framework (Chiu et al., 2022), a state-of-the-art self-supervised learning method for speech representation learning. Best-RQ operates by masking and reconstructing speech units, using a quantization module to discretize continuous representations. This approach enables the model to learn general-purpose speech representations without requiring any labeled data, making it highly scalable to vast amounts of unlabeled audio data. Notably, a key innovation in our implementation of Best-RQ lies in the initialization strategy. Drawing inspiration from our recent findings (An et al., 2024c), where we demonstrate that layers in pre-trained text LLM can effectively initialize the encoder in ASR systems, we initialize the Best-RQ encoder with weights from layers of a pre-trained text LLM—specifically, Qwen3 model (Yang et al., 2025a). This cross-modal initialization strategy is based on the hypothesis that the deep linguistic and semantic knowledge encoded in the LLM can provide a beneficial inductive bias for learning speech representations. We observe that initializing Best-RQ with a pre-trained text LLM significantly accelerates the training convergence and improves the quality of the learned speech representations compared to random initializations.
Stage 2: Supervised pre-training via attention-based encoder-decoder (AED) framework. The second stage involves supervised pre-training of the audio encoder within a conventional attention-based encoder-decoder architecture. This method follows a well-established and empirically validated training paradigm used in our prior works, specifically the SenseVoice-Large model (An et al., 2024a). In this setup, the encoder is trained end-to-end on large-scale labeled ASR datasets, using standard sequence-to-sequence learning objectives. The primary objective of this supervised pre-training phase is to obtain an encoder that has learned rich acoustic and linguistic features from transcribed speech data. Once trained, the encoder of the AED framework is used to initialize the audio encoder in the downstream LLM-ASR system. This initialization provides a strong starting point for subsequent joint training of the audio and language components, reducing the need for extensive low-level feature learning from scratch and thereby accelerating the training convergence.
@@ 4.2 Supervised Fine-tuning
Stage 1: The parameters of the pre-trained audio encoder and the LLM are kept frozen, while the adaptor module is trained to align the audio encoder’s output representations with the LLM’s semantic space. The training data for this stage is about 200k hours
Stage 2: The LLM parameters are still kept frozen, while the audio encoder and the adaptor module are trained to learn a better semantic representations. This stage uses about 10M hours of low-cost ASR training data and trains one epoch.
Stage 3: The encoder and the adaptor module are frozen, while we update the LLM parameters with Low-Rank Adaptation (LoRA). The purpose of LoRA-based LLM adaptation is to preserve the model’s text generation capabilities while ameliorating catastrophic forgetting of the pre-trained knowledge. This LoRA fine-tuning stage uses 20K hours of ASR data
Stage 4: Full-parameter fine-tuning is applied to both the audio encoder and adaptor, while LoRA is employed to fine-tune the LLM simultaneously. In this stage, we only use the high quality data, which contains about 3M hours of speech. The transcriptions are evaluated by three different ASR models, including Whisper-Large-V3, FireRed-ASR, and SenseVoice.
@@ 4.3 Contextual Supervised Fine-tuning
As a content prior, context can effectively help the model identify and disambiguate key text content from easily confusable pronunciation in ASR tasks, and improve the accuracy of long-term continuous recognition in complex scenarios. Consequently, after the SFT training (Section 4.2), we further train FunAudio-ASR on the contextual and long-duration data to expand its contextual modeling capability.
@@
Step 2: Relevant Context Synthesis: We use the Qwen3-32B model to synthesize contextual content. Given the current conversation content and the extracted keywords, we prompt Qwen3-32B to synthesize multiple, diverse contextual content that align with spoken conversation characteristics. For the synthesized context, we then use keyword matching to filter out contextual pieces that do not contain the specified keywords. If no keywords are extracted from the current conversation in the previous step, the LLM is prompted to synthesize context based solely on the current conversation content.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We evaluate the training efficiency of FunRL on 8 A100 GPUs, with results shown in Figure 4(b). For approximately one hour of input audio, one training step takes about 54.6 seconds, yielding a real-time factor (RTF) of approximately 0.015. As shown in Figure 4(b), the SGLang rollout phase dominates the computation time, while device-switching overhead accounts for less than 6% of the total computation time. This indicates that the strategy of alternating GPU utilization in FunRL is highly efficient, making it a scalable and effective solution for RL training in LALMs.
@@
Figure 4: The framework and time consumption analysis for our FunRL.
@@ 4.4.3 Constructing the RL Training Data
Addressing practical issues in application scenarios, we construct a small but high-quality RL training data using the following approach.
  • Hardcase Samples. We collect a large amount of unlabeled speech and transcribe each utterance using the FunAudio-ASR model after contextual SFT along with three other distinct ASR systems including Whisper, FireRed-ASR and SenseVoice. If the outputs from the three external systems are consistent with each other (WER < 5%) but differ significantly from FunAudio-ASR’s result (WER > 10%), the sample is identified as a hard case and included in the RL training set.
  • Long-duration Samples. We select audio segments longer than 20 seconds to improve model performance on extended speech inputs, which are common in real-world applications but leak in the training data (less than 10%).
  • Hallucination-related Samples. We specifically include data where the base model exhibits hallucination behavior, which could be significant longer than the ground-truth or contains repeated phase. Additionally, we incorporate the utterances with the reference transcripts containing long repetitions of words or phrases—cases that resemble hallucinations but are genuinely present in the audio—to help the model distinguish between real and spurious patterns.
@@
The final RL training data consist of 100K samples, with 20K utterances in each of the above five subsets. Thanks to the efficiency of the FunRL framework, the entire RL training of FunAudio-ASR can be completed within one day with 8 A100 GPUs.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Mandarin Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary FunAudio-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual FunAudio-ASR model to support these languages. The current multilingual FunAudio-ASR model (denoted by FunAudio-ASR-ML) supports Mandarin Chinese, English, Vietnamese, Thai, and Indonesian, with a broader language coverage planned for future releases. During training, we downsample the Mandarin and English data used for the Chinese-English FunAudio-ASR model and upsample the Vietnamese, Thai, and Indonesian data to balance the distribution. In total, this multilingual dataset comprises approximately 500K hours of audio. The training methodology is the same as that used for the Chinese-English FunAudio-ASR model (Section 4).

### v2

@@ 1 Introduction
  • Optimization for practical production usage. Beyond achieving state-of-the-art performance on standardized benchmarks, FunAudio-ASR is meticulously engineered to meet the complex demands of real-world deployment scenarios, with a particular focus on practical usability, reliability, and user experience. We implement a comprehensive suite of optimizations in multiple dimensions, each addressing specific challenges encountered in commercial applications. (1) First, we implement a highly efficient streaming ASR architecture for FunAudio-ASR that supports real-time processing with minimal latency, enabling seamless integration into live applications such as video conferencing, live captioning, and voice-controlled devices. (2) Second, we enhance noise robustness substantially through a multi-stage approach. (3) Third, we implement advanced code-switching capabilities that seamlessly handle transitions between Chinese and English within the same utterance, which is critical for multilingual users in global business environments. (4) Fourth, we integrate customizable hotword recognition that allows users to define domain-specific terms or phrases for enhanced recognition accuracy. This feature is particularly valuable in specialized domains such as healthcare, enterprise, and automotive technology.
    To thoroughly evaluate these production-oriented optimizations, we develop a comprehensive evaluation protocol that includes both standardized benchmarks and real-world usage scenarios. This protocol encompasses distinct test sets, each simulating different application contexts. Our evaluation results reveal that FunAudio-ASR not only excels in recognition accuracy but also provides superior practical performance. These results demonstrate that FunAudio-ASR successfully bridges the gap between academic research and commercial production readiness, offering a comprehensive solution to address real-world speech recognition challenges.
@@ 2 Model Architecture
FunAudio-ASR comprises four key components: (1) an audio encoder that extracts representations from the input speech, implemented as multiple layers of transformer encoder, (2) an audio adaptor that connects the audio encoder output with the LLM, implemented as two layers of transformer encoder, (3) a CTC decoder that is built upon the audio encoder to obtain the initial recognition hypothesis, which will be used for hotword customization, as described in Section 5.5, and (4) an LLM-based decoder that produces output based on the audio condition and the CTC prediction.
To address varying constraints on computational resources and requirements on inference efficiency, we propose two models with different model sizes: FunAudio-ASR and FunAudio-ASR-nano. FunAudio-ASR comprises an audio encoder with 0.7B parameters and an LLM-based decoder with 7B parameters, aiming for high recognition accuracy; FunAudio-ASR-nano comprises an audio encoder with 0.2B parameters and an LLM-based decoder with 0.6B parameters, seeking to strike a balance between accuracy and efficiency to meet the demands of low-computational-resource scenarios.
@@ 3.2 Supervised Fine-tuning Data
The supervised fine-tuning (SFT) data consist of approximately millions of hours of data, including: human-transcribed data, pseudo-labeled data, environmental noise data, TTS generated data using CosyVoice3 (Du et al., 2025), simulated streaming data, noise augmented data, and hotword customized data.
@@ 4.1 Pre-training of Audio Encoder
To develop a robust and effective audio encoder for integration into an LLM-based ASR (LLM-ASR) system, we adopt two complementary approaches, as shown in Figure 3. These strategies aim to leverage both self-supervised and supervised learning paradigms to produce high-quality speech representations that can be effectively aligned with the linguistic knowledge in the LLM.
Stage 1: Self-supervised pre-training via Best-RQ framework with pre-trained text LLM initialization. The first stage leverages self-supervised pre-training through the Best-RQ (BERT-based Speech pre-Training with Random-projection Quantizer) framework (Chiu et al., 2022), a state-of-the-art self-supervised learning method for speech representation learning. Best-RQ operates by masking and reconstructing speech units, using a quantization module to discretize continuous representations. This approach enables the model to learn general-purpose speech representations without requiring any labeled data, making it highly scalable to vast amounts of unlabeled audio data. Notably, a key innovation in our implementation of Best-RQ lies in the initialization strategy. Drawing inspiration from our recent findings (An et al., 2024c), where we demonstrate that layers in pre-trained text LLM can effectively initialize the encoder in ASR systems, we initialize the Best-RQ encoder with weights from layers of a pre-trained text LLM—specifically, Qwen3 model (Yang et al., 2025a). This cross-modal initialization strategy is based on the hypothesis that the deep linguistic and semantic knowledge encoded in the LLM can provide a beneficial inductive bias for learning speech representations. We observe that initializing Best-RQ with a pre-trained text LLM significantly accelerates the training convergence and improves the quality of the learned speech representations compared to random initializations.
Stage 2: Supervised pre-training via attention-based encoder-decoder (AED) framework. The second stage involves supervised pre-training of the audio encoder within a conventional attention-based encoder-decoder architecture. This method follows a well-established and empirically validated training paradigm used in our prior works, specifically the SenseVoice-Large model (An et al., 2024a). In this setup, the encoder is trained end-to-end on large-scale labeled ASR datasets, using standard sequence-to-sequence learning objectives. The primary objective of this supervised pre-training phase is to obtain an encoder that learns rich acoustic and linguistic features from transcribed speech data. Once trained, the encoder of the AED framework is used to initialize the audio encoder in the downstream LLM-ASR system. This initialization provides a strong starting point for subsequent joint training of the audio and language components, reducing the need for extensive low-level feature learning from scratch and thereby accelerating the training convergence.
@@ 4.2 Supervised Fine-tuning
Stage 1: The parameters of the pre-trained audio encoder and the LLM are kept frozen, while the adaptor module is trained to align the audio encoder’s output representations with the LLM’s semantic space. The training data for this stage is about 200k hours and this stage takes about 70K training steps.
Stage 2: The LLM parameters are still kept frozen, while the audio encoder and the adaptor module are trained to learn better semantic representations. This stage uses about 10M hours of low-cost ASR training data and trains one epoch.
Stage 3: The encoder and the adaptor module are frozen, while we update the LLM parameters with Low-Rank Adaptation (LoRA). The purpose of LoRA-based LLM adaptation is to preserve the model’s text generation capabilities while ameliorating catastrophic forgetting of the pre-trained knowledge. This LoRA fine-tuning stage uses 20K hours of ASR data.
Stage 4: Full-parameter fine-tuning is applied to both the audio encoder and the adaptor, while LoRA is employed to fine-tune the LLM simultaneously. In this stage, we only use the high quality data, which contains about 3M hours of speech. The transcriptions are evaluated by three different ASR models, including Whisper-Large-V3, FireRed-ASR, and SenseVoice.
@@ 4.3 Contextual Supervised Fine-tuning
As a content prior, context can effectively help the model identify and disambiguate key text content from easily confusable pronunciation in ASR tasks, and improve the accuracy of long-term continuous recognition in complex scenarios. Consequently, after the SFT training (Section 4.2), we further train FunAudio-ASR on the contextual and long-duration data to enhance its contextual modeling capability.
@@
Step 2: Relevant Context Synthesis: We use the Qwen3-32B model to synthesize contextual content. Given the current conversation content and the extracted keywords, we prompt Qwen3-32B to synthesize multiple, diverse contextual content pieces that align with spoken conversation characteristics. For the synthesized context, we then use keyword matching to filter out contextual pieces that do not contain the specified keywords. If no keywords are extracted from the current conversation in the previous step, the LLM is prompted to synthesize context based solely on the current conversation content.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We evaluate the training efficiency of FunRL on 8 A100 GPUs, with results shown in Figure 4(b). For approximately one hour of input audio, one training step takes about 54.6 seconds, yielding a real-time factor (RTF) of approximately 0.015. As shown in Figure 4(b), the SGLang rollout phase dominates the computation time, while device-switching overhead accounts for less than 6% of the total computation time. This observation indicates that the strategy of alternating GPU utilization in FunRL is highly efficient, making FunRL a scalable and effective solution for RL training in LALMs.
@@
Figure 4: The overview and time consumption analysis for our FunRL framework.
@@ 4.4.3 Constructing the RL Training Data
For addressing practical issues in real-world application scenarios, we construct a small but high-quality RL training data using the following approach.
  • Hardcase Samples. We collect a large amount of unlabeled speech and transcribe each utterance using the FunAudio-ASR model after contextual SFT (referred to as the base model), along with three other distinct ASR systems, including Whisper, FireRed-ASR, and SenseVoice. If the outputs from the three external systems are consistent with each other (WER < 5%) but differ significantly from FunAudio-ASR’s output (WER > 10%), the sample is identified as a hard case and is included in the RL training set.
  • Long-duration Samples. We select audio segments longer than 20 seconds to improve model performance on extended speech inputs, which are common in real-world applications but scarce in the training data (less than 10%).
  • Hallucination-related Samples. We specifically include data where the base model exhibits hallucination behavior, which could be significantly longer than the ground truth or contain repeated phases. In addition, we incorporate the utterances with the reference transcripts containing long repetitions of words or phrases, cases that resemble hallucinations but are genuinely present in the audio, to help the model distinguish between real and spurious patterns.
@@
The final RL training data consist of 100K samples, with 20K utterances in each of the five subsets described above. Thanks to the efficiency of the FunRL framework, the entire RL training of FunAudio-ASR can be completed within one day with 8 A100 GPUs.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary FunAudio-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual FunAudio-ASR model to support these languages. The current multilingual FunAudio-ASR model (denoted by FunAudio-ASR-ML) supports Chinese, English, Vietnamese, Thai, and Indonesian, with a broader language coverage planned for future releases. During training, we downsample the Chinese and English data used for the Chinese-English FunAudio-ASR model and upsample the Vietnamese, Thai, and Indonesian data to balance the distribution. In total, this multilingual dataset comprises approximately 500K hours of audio. The training methodology is the same as that used for the Chinese-English FunAudio-ASR model (Section 4).